How do you do this problem with work on how you did it?
Add these complex numbers. Write your in a + bi form.
(-2+3i)+(-1-4i)

Found 2 solutions by jim_thompson5910, Simnepi:
Answer by jim_thompson5910(35256)

(-2+3i)+(-1-4i)

-2+3i-1-4i

(-2-1)+(3i-4i)

(-3)+(-1i)

-3 - i

It's now in a + bi form where a = -3 and b = -1

Answer by Simnepi(216)

Add the real parts and the imaginary parts separately as if they were different terms, like x and y.
So
-2 + -1 = -3 (the real part)
and
(+3i) + (-4i) = -i (the imaginary part)
giving
(-2+3i)+(-1-4i) = (-3-i)